OpenStudy (anonymous):

Find the perimeter of this triangle Below are each sides 5a^2+12a-8 14a^2-8a-9 a-5

OpenStudy (rogue):

Just add up the lengths of each side to find the perimeter.\[(5a^2+12a-8) + (14a^2-8a-9) + (a-5) = 5a^2 + 14a^2 + 12a - 8a + a - 8 - 9 - 5\]\[5a^2 + 14a^2 + 12a - 8a + a - 8 - 9 - 5 = 19a^2 + 5a - 22\]

OpenStudy (anonymous):

P = ( 5a^2+12a-8) + (14a^2-8a-9) +(a -5) = 19a^2 + 5a -22

OpenStudy (anonymous):

I noticed some signs were changed? My answer is 1/38....That can't be right?? Can it?

OpenStudy (phi):

both answers above are the same (and correct). How did you get a number. Do you have a value for a?

OpenStudy (anonymous):

I solved 19a^2+5a-22

OpenStudy (phi):

tol solve it, you need an equation. For example 19a^2+5a-22=0 (which can't be the case here, if it really is the perimeter, it is not zero.) unless you have left out some info, the problem just give us expressions that we add up.